OpenStudy (anonymous):

charles opened the trunk and found $6750 in $1 bills and $10 bills. If there were 150 more ones than tens, how many of each kind were there?

OpenStudy (anonymous):

Let\[n_1, n_{10}\]be the number of $1 and $10 bills respectively. You know that\[n_1=n_{10}+150\]since the number of 1s is 150 more than the number of 10s. You also know\[6750=n_1$1+n_{10}$10\]that is\[6750=n_1+10n_{10}\]

OpenStudy (anonymous):

Sub. the expression for n_1 into the second equation to find\[6750=(n_{10}+150)+10n_{10}=11n_{10}+150\]

OpenStudy (anonymous):

Then\[n_{10}=\frac{6750-150}{11}=600\]

OpenStudy (anonymous):

There are 600 $10 bills. So the number of $1s is,\[n_1=6750-10 \times 600=750\]

OpenStudy (anonymous):

thanks so very much

OpenStudy (anonymous):

No probs :)